Chuck Conlon is the materials manager at Spencer Electric in Muncie, Indiana. Chuck is looking for ways to reduce inventories. A recent statement shows the following inventories: raw materials, $5,127,400; work-in-process, $6,432,000; and finished goods, $3,587,200. This year’s cost of goods sold will be about $34.7 million. Assume 50 business weeks per year.

A. What is the raw material inventory         

B. What is the finished goods inventory         

C. What is work in process inventory                                         

D. Average aggregate inventory value is  

E. Weekly sales is

1F. Express total inventory as weeks of supply

  1G. Express total inventory as Inventory turns. (inventory turnover                                                                                                                                  
Answer #1

A. What is the raw material inventory   

raw material inventory = $5,127,400

B. What is the finished goods inventory

finished goods inventory = $3,587,200   

C. What is work in process inventory

work in process inventory = $6,432,000   

D. Average aggregate inventory value is

$5,127,400+$3,587,200+$6,432,000 = 15146600

E. Weekly sales is (34.7*10^6)/50 = 694000

1F. Express total inventory as weeks of supply

Weeks of supply = Average aggregate inventory value/weekly sales (at cost) = 15146600/694000 = 21.82507205

1G. Express total inventory as Inventory turns.
Inventory turns = Annual sales at cost/Average aggregate inventory value = (34.7*10^6)/15146600 = 2.290943182